ITES2001 Fourth Italian-Spanish conference on General Topology and its applications Bressanone, 27-30 June 2001

Metric shape and P -embeddings

Luciano Stramaccia Università di Perugia, Italy

We use some techniques from (metric) shape theory to give new insights on the Borsuk’s characterization theorems for AR and ANR spaces and the homotopy extension theorem.